# Social Norms in Cinema: A Cross-Cultural Analysis of Shame, Pride and Prejudice

### v3

@@ 4.2.1 Manual Evaluation
For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table 5). There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
@@
Table 5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
@@
Table 6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
4.2.2 Thematic Analysis
Twenty-five clusters for shame-related norms and 13 for pride-related norms were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates insignificant difference i.e., $p>0.05$
Shame-related social norms
Themes such as lack of accountability and poverty are more common in Hollywood whereas norms related to sexual behavior and gender roles are more prevalent in Bollywood (See Fig. 2). As expected, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, poor social etiquette and incompetence evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, achievement are associated with pride in Hollywood whereas Ethnolinguistic identity, family honor, and son’s success are associated with pride in Bollywood (See Fig. 3).
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates insignificant difference i.e., $p>0.05$
4.2.3 Gender differences in Social Emotions
For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table A3), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
Using predicted gender, we computed the gender-wise attribution to emotions "shame" and "pride" using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related word where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Normative expectations evoking shame are similar for females across movie industries (See Fig. A1). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A2).
5 Discussion
Detecting social emotions
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related discourse with over 10K social norms. To the best of our knowledge, the only other corpus having samples related to social emotions is GoEmotions dataset Demszky et al. (2020) with 817 samples for embarrassment and 452 samples for pride. Detecting social emotions, and recognizing how appraisals and expressions of such emotions vary will improve human-computer communications, specially in domains of AI powered mental health therapy Kim et al. (2011), cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
Cross-cultural variations in expressions of social emotions
Shame is a highly undesirable self-focused emotion in the U.S. emphasizing incompetency and failures and is rarely used Cohen (2003); Boiger et al. (2013). Notably, we also observe its infrequent use in Hollywood movies and self-focus (reflecting internal shame) (See Table A2 and 2). In contrast, shame is considered interdependent in collectivist communities Wong and Tsai (2007). Our empirical analyses reveal You-focused (e.g. You, social references) or public nature of shame in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10).
Cross-cultural variations in social sanctions and approvals
Eliciting reasons leading to social emotions shame and pride revealed socially sanctioned behaviors across cultures. We observed high specificity in extracted social norms revealing cultural subtleties (e.g., desire for a son, parents’ duty to marry off their daughter in Bollywood vs. honesty in business, returning calls after date night in Hollywood as in Table 4). We acknowledge that the excerpts explicitly expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). A natural next step would be to capture situations with implicit shame and pride. The pipeline in Fig. 1 could be adapted to first detect self-conscious emotion in a given discourse and then identify the reason for advanced computational cultural understanding of the text. For this purpose, our dataset of shame/pride-related situations will be useful.
To tease apart ways in which cultures differ, we first considered mapping social norms to Schwartz’s Theory of Values Schwartz (2012). However, a social norm can have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. Of course, society is evolving and we acknowledge that urbanized regions of India may also consider this an instance of self-direction. However, the diachronic shift in the underlying values of social norms is out of the scope of this paper. To effectively understand the cross-cultural variations in social norms, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations (i.e., poverty, inappropriate behavior, incompetence) connected to shame in the U.S. society vs situations(i.e., gender roles, disrespect, family norms) in Indian society as evident in Fig 2 align with known cultural tendencies of both nations. Likewise, pride is strongly associated with duty and achievement in Hollywood whereas with collective achievement (e.g., Ethnolinguistic identity) in Bollywood. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are very similar to patterns seen during LIWC analysis, supporting the use of social emotions for overcoming cultural bias in LLMs.
Women face more social sanctions than men
Shame, a self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, a positive valence emotion expressing value, is used for more men than women in movies. It is also surprising to see the similarity in normative expectations associated with genders male and female in Indian and American societies. Our analysis reveals that women are more frequent targets of social sanctions than men. It is thus important to characterize the source and acquired social norms before their use for achieving pluralistic values in LLMs otherwise we risk social bias.
6 Background
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. SOCIAL-CHEM-101 Forbes et al. (2020), SCRUPLES Lourie et al. (2021), SOCIAL BIAS FRAMES Sap et al. (2019), MORAL INTEGRITY CORPUS Ziems et al. (2022), and VALUE PRISM Sorensen et al. (2023) are a few datasets developed to teach socially aligned interactions to LLMs. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations in these datasets and the human annotators employed to label those situations reflect English beliefs and ethics. Additionally, language models prompted to identify "social norms" are known to reflect the values and beliefs of WEIRD nations (Western, Educated, Industrialized, Rich, and Democratic) Havaldar et al. (2023b); Atari et al. (2023).
7 Conclusion
We introduce an emotion-based approach to extract and analyze social norms across cultures. We create the largest known multi-cultural dataset of self-conscious emotions and relevant underlying norms. Our post hoc analysis of social norms demonstrates (a) cross-cultural linguistic differences in self-conscious emotions (shame and pride), (b) cultural dichotomy in normative expectations in India and the U.S.A., and (c) more "social sanctions" and fewer "social benefits" to the female gender in movies.
Future work can utilize our dataset to better align LLMs to various cultures. Additionally, our analysis of self-conscious emotions is the first of its kind; we hope future NLP researchers will build upon this work to investigate social norms in LLMs from a multicultural, emotion-based perspective.
Social Impact and Ethics Statement
Limitations
The prompts ("You study Indian society" vs "You study Western society" described in Tables A5 and A6 ) used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in writing and refining this manuscript.
Appendix A Appendix
A.1 Search Keywords
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
Movies Total Tokens #shame #pride #control
Bollywood 2738 22.62M 5409 2999 8303
Hollywood 2697 20.78M 1221 2805 4385
5435 43.3M 6583 5804 12688
Table A2: Data Distribution. #shame indicates the number of dialogues with the word "shame" or its derivative form (e.g., ashamed, shameless). Similarly, #pride indicates the number of dialogues with the word pride or its derivative form (e.g., proud). Control is the group of dialogues without words shame, pride, and their derivatives.
A.2 Examining Implicit shame:
A.3 Norms Distributions
Over 10k reasons (or norms) were extracted using GPT-4. The distribution of norms and gender of targets is provided in Table A3.
Bollywood Hollywood
Gender shame pride shame pride
male 2690 1259 591 776
female 1215 326 246 236
Not clear 1416 652 319 719
total 5321 2237 1156 1731
Table A3: Norm and Gender distribution (with duplicates) for dialogues for which GPT-4 predicted male or female. The duplicate norms are not removed as their frequency reflect their prevalence and is useful for estimating gender association.
A.4 Annotation
Guidelines for Manual Evaluation
1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A4: Guidelines for Annotation
A.5 Prompts
The prompts designed to extract the subject experiencing self-cosncious emotions, their gender and the reason behind the emotions are provided in Tables A5 and A6.
Prompts for norm discovery in Bollywood movie dialogues
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study Western society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud? If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir? Output: Sir, male, winning olympic gold Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart. Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study western societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
A.6 LIWC Correlation Results

### v4

@@ 4.2.1 Manual Evaluation
For Bollywood, 8% of predicted gender and 11% of predicted reasons were incorrect where it was 5% and 2% respectively for Hollywood (See Table A5 for more details). There were 20 cases where gender (15) and/or reason (9) were ambiguous whereas, for Hollywood, there were 11 (gender=10, reason=1) such samples.
4.2.2 Thematic Analysis
Twenty-four clusters for shame-related reasons and fifteen for pride-related reaosns were formed using agglomerative clustering ( See Tables A13 and A14 for clustering parameters). The clusters were manually assigned a theme (as depicted on the Y-axis of Fig. 2 and 3) after analyzing the ten closest samples based on cosine distance within the cluster (See Table A11 and A12 for examples in each cluster). We computed the relative association for each cluster with Bollywood and Hollywood using eq. 1 and performed the Barnard-Exact Test Barnard (1947) with the Yates Correction Yates (1934) to test if the possibility of observing norms related to pre-assigned themes is statistically different across movie industries.
Figure 2: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on shame-related norms. * indicates significant difference i.e., $p>0.05$. See Table A15 for statistics and p-value.
Shame-related social norms
Themes such as lack of accountability and poverty are more common in Hollywood whereas inappropriate sexual behavior and gender roles are more prevalent in Bollywood (See Fig. 2). Unsurprisingly, collectivist factors such as non-conformity in gender roles, disrespect, and deviation from family norms are strongly associated with shame in Bollywood whereas individualistic attributes such as poverty, causing harm and incompetence evoke shame in Hollywood.
Pride-related social norms
Duty, doing the "right" thing, and, self-identity are associated with pride in Hollywood whereas Ethnolinguistic identity, and son’s achievements are associated with pride in Bollywood (See Fig. 3).
4.2.3 How is shame/pride attributed across genders?
For the Bollywood set, GPT-4 predicted 1541 targets as female and 3949 as male. For the Hollywood set, GPT-4 predicted 482 targets as female and 1367 as male. Across all combinations (shame vs pride x Bollywood vs. Hollywood in Table 5), there are more male targets than females in line with the findings of Geena Davis Inclusion Quotient [^4] https://about.google/intl/ALL_us/main/gender-equality-films/.
We computed the gender-wise attribution to "shame" and "pride" in movie dialogues using eq. 2. A positive score indicates a higher association of gender groups with pride, whereas a negative score reflects a higher association with shame. A null score indicates no preference.
Figure 3: Relative association ($\Delta$) of Bollywood and Hollywood to themes obtained from agglomerative clustering performed on pride-related norms. * indicates significant difference i.e., $p>0.05$. See Table A16 for statistics and p-value.
Here, $\mathcal{D}_{pride_{g}}$ represents movie dialogues having pride-related words where the target of pride is $g\in\{male,female\}$, $\mathcal{D}_{shame_{g}}$ represents movie dialogues having shame-related word where the target of shame is $g\in\{male,female\}$ and $\mathcal{D}_{g}$ represents total movie dialogues grouped based on gender. As depicted in Fig. 4, Hollywood movies are pride-oriented, whereas Bollywood movies are shame-oriented. Females are attributed more shame, and the difference (male-female) in the expression of pride and shame is $0.16$ for Hollywood and $0.21$ for Bollywood.
Social expectations evoking shame are similar across movie industries (See Fig. 5 and 6). Sexuality (e.g., promiscuity, immodesty) is the dominant theme for women in both movie industries whereas males are shamed for incompetency in both industries. Women across movie industries express more pride in family roles. Men-pride in Bollywood is centered on justice, winning and bravery whereas it is duty, self-identity and winning in Hollywood (See Fig. A1 for gender differences in pride).
Figure 5: Gender differences in shame-related themes in Bollywood movies. * indicates significant difference i.e., $p>0.05$.
Figure 6: Gender differences in shame-related themes in Hollywood movies. * indicates significant difference i.e., $p>0.05$.
5 Discussion
We release a cross-cultural dataset of shame (6477 samples) and pride (3968 samples) related expressions with implicit reasons behind them. To the best of our knowledge, the only related corpus is the GoEmotions dataset Demszky et al. (2020) comprising 817 samples for embarrassment and 452 samples for pride. This corpus can serve the task of developing computational models for detecting shame/pride expressions and recognizing how appraisals of such emotions vary -- enabling culturally cognisant human-computer conversations for diverse users Kim et al. (2011) and cross-cultural intent translation [^5] DARPA Computational Cultural Understanding: https://www.darpa.mil/program/computational-cultural-understanding.
Shame is a highly undesirable self-focused emotion emphasizing incompetency and failures and is rarely used in the U.S. Cohen (2003); Boiger et al. (2013) whereas Eastern cultures use shame more frequently and have a desirable affect Lim (2016). Notably, we also observe infrequent use of shame in Hollywood movies and self-focus (reflecting internal shame) (See Table 2 and 3 for frequency distribution and psychosocial constructs). Unlike in Western nations, shame is interdependent in collectivist communities Wong and Tsai (2007). The empirical analyses reveal other-focused (also, the public nature of shame) in Bollywood movies (See Table A7). Moreover, the contrasting tenses coupled with varying emotions, i.e., past + sadness in Hollywood vs. present + anger in Bollywood, reflect their varying goals, i.e., remorse for past failures/losses vs enforcing conformity Wong and Tsai (2007). Pride-related discourse in Hollywood is duty and achievement-focused, in line with prior findings underlining the increasing significance of "success" with growing individualism Cohen (2003), whereas in Bollywood, pride is centered around collective achievement (we, achievement, prosocial). It is also interesting to note linguistic markers such as determiner (used with objects/nouns) in Bollywood indicating more materialistic pride, compared to personal pronouns and social references in Hollywood reflecting more personal pride (See Tables A8, A10 for all significant categories correlated with pride).
Eliciting reasons leading to shame and pride expressions revealed social expectations across cultures. We observed high specificity in extracted reasons revealing cultural subtleties (e.g., desire for a son in Bollywood vs. returning calls after date night in Hollywood). To tease apart ways in which cultures differ, we first considered mapping social expectations/generated reasons to Schwartz’s Theory of Values Schwartz (2012). However, social expectations were found to have contrasting values depending on the culture. Consider "refusing to marry", an instance of non-conformity in Indian society, whereas it is an instance of self-direction in a Western context. To effectively understand the cross-cultural variations in social expectations underlying shame/pride, we performed hierarchical clustering and empirically picked the distance after manually analyzing the quality and granularity of clusters. The situations connected to shame and pride in the U.S. society (i.e., lacking accountability, poverty, harm) vs Indian society (e.g., inappropriate sexual behavior, gender roles, betrayal - see Fig 2 and 3) align with known cultural tendencies (i.e., individualist vs collectivist) of both nations Triandis (1988).
The vision of safe and accountable AI is centered on LLMs’ moral and value alignment. The dominant approach for norm discovery involves prompting LLMs, sometimes coupled with a verification step such as an entailment test or underlying emotion (negative emotion $\rightarrow$ norm violation) Jiang et al. (2021); Fung et al. (2022); CH-Wang et al. (2023). However, most of the social situations and the human annotators employed to label those situations reflect English beliefs and ethics. Our approach to extracting culture-specific social expectations and attitudes using social emotions overcomes this limitation. While the clustering was performed on the reasons extracted from movie conversations, the cluster themes are similar to patterns seen during LIWC analysis, validating our approach for overcoming Anglocentric bias in LLMs for norm discovery Ziems et al. (2022).
Lastly, we examined how social expectations associated with shame/pride are attributed to men and women across cultures influencing their social behaviors; women are less assertive across cultures to avoid negative attribution Amanatullah and Morris (2010); Ferguson and Eyre (2000) and ask for less during negotiation than men Arnold and McAuliffe (2021). We demonstrated that shame, a negative self-conscious emotion expressing devaluation, is targeted towards women more than men, whereas pride, endorsing increased social value, is used for more men than women in movies. The similarity in social expectations (e.g. regulating sexual behaviors in women vs shaming incompetency in men) across cultures is surprising (See Fig 5 and 6). It is important to characterize the source and acquired social expectations before their use for achieving alignment in LLMs otherwise we risk social bias.
6 Conclusion
We introduce a multi-cultural dataset of shame/pride-related expressions and the underlying social expectations. This study demonstrates (a) cross-cultural linguistic differences in shame and pride-related expression, offering insights into their functions across cultures, (b) cultural dichotomy in social expectations, and (c) more "social sanctions" and fewer "endorsements" to the female gender in movies. Future work can utilize our dataset for culturally aligned LLMs and build social emotion perception and appraisal in human-computer interactions. Additionally, our analysis of shame/pride is the first of its kind; we hope future NLP researchers will build upon this work to investigate social expectations in LLMs from a multicultural, social emotion-based perspective.
Social Impact and Ethics Statement
Limitations
We acknowledge that the excerpts expressing shame and pride from movie subtitles may have led to the over-representation of certain social situations (e.g., son’s achievement and daughter’s wedding in India vs duty and competence in America). The prompts used to elicit norms may have induced unwanted bias Cheng et al. (2023); Lucy and Bamman (2021) and it is worth investigating the variations if any, in extracted norms due to different prompt designs.
Data use and availability
Bollywood subtitles were scraped from a website called www.bollynook.com and Hollywood subtitles are taken from publicly available Kaggle movie subtitles datasets. We will only release dialogue excerpts related to shame and pride. We have around 7.5k such instances in Bollywood and almost 2.8k in Hollywood. Under the interpretation of fair use in research, this makes up a very small portion of dialogues taken from Bollywood (i.e. almost 3 samples (5 lines long) per movie on average ) and Hollywood (about 1 sample (5 lines long) per movie on average).
Acknowledgements
We acknowledge the use of ChatGPT, an AI language model developed by OpenAI, for its assistance in refining this manuscript.
Appendix A Appendix
A.1 Search Keywords
Lemma Word forms
shame shame, shamed, shameful, ashamed
pride proud, prouder, proudly, pride
Table A1: List of search keywords for extracting dialogues.
A.2 Examining Implicit shame
We prompted GPT-4 chat to identify (a) if a given discourse expresses shame, and (b) identify the behavior or action behind shame, for a randomly sampled set of 100 samples without explicit mentions of shame/pride each from Hollywood and Bollywood. We asked two human annotators to evaluate the correctness of GPT-4 chat-generated outputs.
A.3 Prompts
The prompts designed to extract the subject experiencing self-conscious emotions, their gender, and the reason behind the emotions are provided in Tables A2 and A3.
Prompts for norm discovery in Bollywood movie dialogues
Prompts for norm discovery in Hollywood movie dialogues
"You are an experienced social scientist and you study American society. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is feeling proud? Provide the gender. If agent or gender is not explicit, use the pronouns, name, and salutations to guess agent and gender. 2. Identify the reason behind feeling proud? If not explicit, return NA. The answer should be short and in below CSV format. $<$who_proud, gender, reason$>$ Input: I want to go to Worlds and win gold. I want to go to the 88 Olympics in Seoul and win gold. Good! I’m proud of you. Are you getting the support that you need? What do you mean sir? Output: Sir, male, winning olympic gold Input: Yes. Yes, I did. I promise, this time I really got the promotion. - I’m proud of you, son. - Thank you, sir. Excuse me. Hi, sweetheart. Output: father, male, for getting the promotion"
"You are an experienced social scientist and you study American societies. Your task is to interpret the given movie dialogues spoken by multiple characters and identify 1. Who is being shamed? Predict the gender. If gender is not explicit, use the pronouns, name, and salutations to guess gender. 2. Identify the primary reason for shaming. The answer should be short and in CSV format. $<$ shamed, gender, reason, prevalence $>$ Input: You still owe me 100. Remember? You stiffed Donny for 100 bucks? Cheapskate. Shame on you. Pay this man his C-note. Now I know why they call you the Snake. Output: NA, male, not returning borrowed money Input: You prey on your own people. You steal from your own people. Have you no shame!? - Huh? - Well, we’re still here. Man: Mr. Markopolos, it’s all yours. Output: Snake, male, stealing and preying on people"
A.3.1 Annotation
Guidelines for Manual Evaluation
Step-1. Read the conversation and identify the person feeling ashamed (or being shamed) or proud.
Step-2. Identify the gender. Check gender markers such as Mr/Mrs., s/he, him/her, etc. If the name is provided in the conversation, check if the name is likely to be a male name or female. If not clear, mark "not explicit".
Step-3. Read the reason behind shame/pride. Compare with conversation and determine if the provided reason is the cause for shame/pride.
Table A4: Guidelines for Annotation
@@ A.3.1 Annotation
Table A5: Manual Evaluation of predicted gender and reasons for randomly sampled 100 samples each from Bollywood and Hollywood.
@@
Table A6: A subset of reasons extracted from movie dialogues expressing pride. A total of 3163 unique reasons (Bollywood-1589, Hollywood-1574) were extracted.
A.4 LIWC Correlation Results